# A High Efficient and Scalable Obstacle-Avoiding VLSI Global Routing Flow

### v2

@@ 5.1 Initial Routing
The first step of initial routing is to generate OARSMTs for all nets. To reduce runtime while maintaining quality, the OARSMT algorithm only considers obstacles that overlap with the bounding box of each net as part of its inputs. This may result in some nets being violated, but these issues can be quickly addressed in subsequent steps. After generating the OARSMTs for all nets, traditional pattern routing is applied to determine the routing patterns and the layers of the edges. An example of pattern routing with layer assignment is illustrated in Figure 6(b) and (c), which also presents the initial routing solution.
Figure 6: (a) An example of a simple case corresponds to both traditional initial routing and our proposed initial routing. The different colors represent different nets. (b) An example of the 3D result obtained under our initial routing. (c) The corresponding 2D solution of (b).
Note that although the proposed approach in Figure 6(a) places more demands and bends on GCells near obstacles than the traditional one due to the avoiding of obstacles by OARSMTs, it does not infer that the proposed approach incurs greater congestion, as the traditional approach does not even begin to effectively avoid obstacles at this stage. Actually, the proposed approach has even less congestion in the final results, because most of the obstacles are already well avoided at the very early stage, which alleviates the tasks during rip-up and reroute.
@@ 5.2 Rip-up and Reroute
Although initial routing can help most nets avoid obstacles, further attention is needed to address congestion. Therefore, maze routing is applied to rip up and reroute the remaining overflowed and violated nets. However, maze routing is time-consuming due to the extensive search space involved. One relatively effective method to reduce the search space for maze routing, without significantly compromising solution quality, is to use sparse graphs. Unfortunately, in many cases, the paths generated on a sparse graph may not provide obstacles-avoiding solutions, as illustrated in Figure 7(a) and (d). To remedy this, we propose OARSMT-guided sparse maze routing and an enhanced obstacle-aware sparse maze routing approach, both of which significantly improve the likelihood of avoiding obstacles while maintaining limited runtime overhead.
@@
(1) OARSMT-Guided Sparse Maze Routing: In order to significantly reduce overflows while eliminating obstacle violations during the rip-up and reroute phase, we propose OARSMT-guided sparse maze routing in the first stage of rip-up and reroute. During the process of rip-up and reroute for each net, additional sparse grid lines are created by utilizing the OARSMT structure established during the initial routing. These grid lines extend from each edge of the OARSMT and have a specific width. An example is shown as blue lines in Figure 7(b), which follow the extension of the dotted lines. By combining the ordinary sparse graph and the additional grid lines, the OARSMT-guided sparse graph is generated. An example is shown in Figure 7(c). When performing maze routing on the generated graph, a valid solution without any obstacle violations can be obtained, as shown by the solid line example in Figure 7(c). Note that the solution on the OARSMT-guided sparse graph might not be able to be obtained on an ordinary sparse graph such as Figure 7(a). Besides, each net’s additional grid lines are only applied for its own maze routing, so each OARSMT-guided sparse graph has a limited complexity increase than the ordinary one. As a result, the proposed method has comparable execution time with better quality than using the ordinary graph.
(2) Obstacle-aware Sparse Maze Routing: After utilizing OARSMT-guided sparse maze routing, the overflow and obstacle-violated nets experience a significant reduction; however, there are still some nets with obstacle violations. As shown by the dotted OARSMT in Figure 7(e), the OARSMT algorithm considers only obstacles within the bounding box during the tree generation stage, which may result in invalid solutions. To fix this problem, the obstacle-aware sparse maze routing can be applied exclusively to the remaining nets with violations. Utilizing the additional grid lines extending from the boundaries of all obstacles and pins for maze routing, by combining these lines with the ordinary sparse graph, obstacle-aware sparse graph can be created, as shown in Figure 7(e) and (f).The solid lines illustrate a solution for obstacle-aware sparse maze routing. This method ensures that most nets successfully navigate around obstacles. Although the presence of numerous obstacles can slow down the obstacle-aware sparse maze routing process, it is only applied to a limited number of remaining nets. As a result, the overall runtime for this step remains reasonable.
@@ 6.1 Experiment Setup
The experiments are conducted on a server with Intel Xeon Gold 6348 CPU. In the experiments of the OARSMT algorithm, there are two types of benchmarks: Randomized testcases and standard benchmarks. Each testcase in benchmarks contains one net to be test, which follows the regular evaluation process as previous OARSMT works [11, 4, 21]. For randomized testcase generation, 3 factors are set: The number of pins and obstacles, and the obstacle density on the layout. Other parameters are randomly generated. Each setting is tested 50 times to get the average result. Besides, we also test all the standard benchmarks IND, RC, and RT [22] involved in the compared works [11, 4, 21]. $k_{l}=5$ and $k_{m}=2$ in the experiments. For OARSMT comparison, the advanced work proposed by Lin et al. [11], a recent work [21] and a traditional work [4] are involved, with works [11, 4] reproduced. Similar to recent work [12], our experiments also apply the algorithm in work [11] by 2D mode. The 2D mode of work [11] does not trim any key techniques in the 3D mode, and thus it still maintains the quality with faster speed. Besides, as the tree generation in regular global routing flow typically generate 2D trees, applying work [11] under 2D mode is able to perform the comparison under a complete global routing flow.
In the experiments of obstacle-avoiding global routing, the framework of global router CUGR2.0 [20] is adopted, where the proposed algorithms are implemented. We utilize the same metrics as applied in the ICCAD19 global routing competition [23], including wirelength, via count, congestion overflow cost and so on. Meanwhile, we also evaluate the obstacle violations, following the same calculation method as the overflow cost in the competition. ISPD24 contest [24] benchmark circuits were used in our experiments as it contains testcases from the regular size to the very large size. We set the macros in the benchmarks as obstacles for our experiments, of which the areas are forbidden to be crossed by wires at all layers. Besides, we also add 60 random obstacles on each testcase of the benchmarks as additional testcases.
@@ 6.2 OARSMT Comparisons on Randomized Testcases
Table 1: The comparisons with the advanced previous work [11] on the randomized testcases.
@@
Wirelength Difference ($<$0 Stands for Improvements)
@@
The OARSMT algorithm is first tested and compared. During the OARSMT algorithm comparisons, all obstacles in the layout are considered, rather than only considering those in the net’s bounding box as the initial routing flow. In this case, there is no violation in the experiments of OARSMT algorithm itself.
The wirelength difference and speedup of the proposed algorithm is shown in Table 1, compared with advanced work [11]. In Table 1, each row stands for a setting of the number of pins, and each column stands for a setting of the number of obstacles. Each group contains 4 runtime results corresponding to different obstacle densities. One can notice that the proposed algorithm has $\sim$10x - 2700x speedup from small cases to large cases.
The absolute values of the runtime for the proposed algorithm are also tested, and are 0.63ms for small cases and only 0.19s for large cases. This not only shows the fast speed, but also proves the scalability.
Besides, Table 1 also shows the wirelength difference, and negative results mean that the proposed algorithm has a smaller wirelength. The wirelength difference ranges from --5.76% to 2.68%. It is shown that the proposed algorithm can also outperform work [11] in the wirelength for various cases, and the wirelength improvement is 1.14% on average. For relatively large designs, the proposed algorithm shows $>$3% wirelength improvement. The experiments indicate that the proposed algorithm has better wirelength for randomized testcases compared with the previous work, but also spends orders of shorter runtime.
@@
Moreover, we evaluate the wirelength improvement of the enhanced rules as shown in Figure 8, which shows the differences between the runs with and without using the enhanced rules. The algorithm with rule enhancements shows an average -1.88% wirelength improvement. For the cases with fewer pins ($\leq$30) or a large number of obstacles ($\geq$500), when obstacles each edge passes become complicated, it shows a bigger -2.22% and -2.69% wirelength improvement, respectively. The results prove that the rule enhancements, by exploring more design space, provide more opportunities to avoid local optimal solutions and achieve better wirelength.
Finally, we briefly report the wirelength and runtime comparisons with previous 2D OARSMT work [4]. The speedup of the proposed work compared to work [4] ranges from $\sim$11x - 7374x, with 1306x on average. Meanwhile, the wirelength difference ranges from -10.04% to -0.47%. The average wirelength improvement among all cases is 5.03%. The previous work [4] has a worse wirelength because it trims some optimized steiner nodes when building an initial tree structure at the early stages. This again indicates the advantages of the proposed algorithm on both quality and speed.
@@ 6.3 OARSMT Comparisons on Standard Benchmarks
Test cases #Pin/#Obs Wirelength Runtime (s)
@@
For OARSMT algorithms, besides the throughout randomized testcases and comparisons, we also compare the proposed algorithm on all standard benchmarks [22] adopted in the compared works, including the advanced work proposed by Lin et al. [11], the recent work GSLS [21], and the traditional work [4]. Each testcase is to generate one OARSMT for one net within a layout with obstacles. The results are shown in Table 6.3. The speedup of our work ranges from $\sim$150x to 5800x, with trading limited wirelength overhead compared with the works [11, 21], and with even better wirelength compared with the work [4]. Note that for Lin et al. [11], there is no result for RC12 in neither the reproduction nor its report, due to the large runtime. In contrast, the proposed algorithm only needs $\sim$0.2s for RC12. Besides, the runtime of the proposed algorithm also shows the scalability as it increases almost linearly with increasing the problem size, and this is in contrast with recent work GSLS [21]. In this case, with large amount of nets in complete designs, previous work can spend more time for tree generation than a traditional routing flow. In conclusion, the experiments on standard benchmarks also prove the high efficiency and scalability of the proposed algorithm. This is not only because of the theoretical lower time complexity, but also because of the simpler processing.
@@ 6.4 Proposed Obstacle-Avoiding Global Routing Evaluations
As shown in Table 3, we compare the proposed global router (“Ours”) with CUGR 2.0 [20], enhanced CUGR 2.0, and our global routing flow integrated with the advanced OARSMT [11] method (“Lin et al.”). The enhanced CUGR 2.0 is CUGR 2.0 with the increased edge cost for the obstacles during maze routing. The cost is large enough that can effectively make the maze routing avoid the obstacles in the sparse graphs if possible. We separately compare wirelength, via count, obstacle violation cost, congestion overflow cost and runtime. Note that the overflow cost does not include the obstacle violation cost.
In Table 3, CUGR 2.0 performs poorly in terms of obstacle violations. Although the enhanced CUGR 2.0 has shown significant improvement, it still fails to effectively eliminate obstacle violations. In contrast, the proposed global routing flow (“Ours”) can reduce the final obstacle violation to zero, which proves its effectiveness. Additionally, “Ours” demonstrates reductions in wirelength and overflow in most test cases when compared to both the original and enhanced CUGR 2.0, achieving an average improvement of 1.96% in wirelength and a 28.06% reduction in overflow cost to enhanced CUGR 2.0, although with acceptable overhead in terms of count and runtime, as more detours are required to navigate around obstacles. The effectiveness of obstacle elimination and overflow is primarily due to the challenges posed by complex obstacle distributions. In traditional algorithms, treating obstacles similarly to overflow complicates the balancing of costs in each stage. This disrupts the collaboration of stages in a traditional global routing process. Meanwhile, maze routing sparse grids in traditional global routers without OARSMT guidance also limit the opportunities of avoiding obstacles. In this case, addressing obstacle violations directly in the early stage and modifying the later stages to incorporate early guidance allows global routing stages to effectively manage obstacles. This is also supported by the insights presented in Section 6.5.
Besides, to further evaluate the effectiveness of the proposed complete flow, we also replace the OARSMT algorithm in the proposed flow with the advanced OARSMT [11] (“Lin et al.” in Table 3). It shows that although both “Ours” and “Lin et al.” eliminate all obstacle violations in the benchmark, “Lin et al.” spends 9.21x runtime on average than “Ours”, with similar wirelength, via count, and overflow cost. This further demonstrates the efficiency of our obstacle-avoiding global routing flow with the proposed OARSMT algorithm.
@@
  • • The Lin et al. is our proposed obstacle-avoiding global routing flow with the advanced OARSMT [11].

### v3

@@ 5.1 Initial Routing
The first step of initial routing is to generate OARSMTs for all nets. To reduce runtime while maintaining quality, the OARSMT algorithm only considers obstacles that overlap with the bounding box of each net as part of its inputs. This may result in some nets being in violation, but these issues can be quickly addressed in subsequent steps. After generating the OARSMTs for all nets, traditional pattern routing is applied to determine the routing patterns and the layers of the edges. An example of pattern routing with layer assignment is illustrated in Figure 6(b) and (c), which also present the initial routing solution.
Figure 6: (a) An example of a simple case corresponds to both traditional initial routing and our proposed initial routing. The different colors represent different nets. (b) The corresponding 2D solution of (c). (c) An example of the 3D result obtained under our initial routing.
Note that although the proposed approach in Figure 6(a) places more demands and bends on GCells near obstacles than the traditional one due to avoiding obstacles by OARSMTs, it does not imply that the proposed approach incurs greater congestion, as the traditional approach does not even begin to effectively avoid obstacles at this stage. Actually, the proposed approach has even less congestion in the final results, because most of the obstacles are already well avoided at a very early stage, which alleviates the tasks during rip-up and reroute.
@@ 5.2 Rip-up and Reroute
Although initial routing can help most nets avoid obstacles, further attention is needed to address congestion. Therefore, maze routing is applied to rip up and reroute the remaining overflowed and violated nets. However, maze routing is time-consuming due to the extensive search space involved. One relatively effective method to reduce the search space for maze routing, without significantly compromising the solution quality, is to use sparse graphs. Unfortunately, in many cases, the paths generated on a sparse graph may not provide obstacles-avoiding solutions, as illustrated in Figure 7(a) and (d). To remedy this, we propose OARSMT-guided sparse maze routing and an enhanced obstacle-aware sparse maze routing approach, both of which significantly improve the likelihood of avoiding obstacles while maintaining limited runtime overhead.
@@
(1) OARSMT-Guided Sparse Maze Routing: In order to significantly reduce overflows while eliminating obstacle violations during the rip-up and reroute phase, we propose OARSMT-guided sparse maze routing in the first stage of rip-up and reroute. During the process of rip-up and reroute for each net, additional sparse grid lines are created by utilizing the OARSMT structure established during the initial routing. These grid lines extend from each edge of the OARSMT and have a specific width. An example is shown with blue lines in Figure 7(b), which follow the extension of the dotted lines. By combining the ordinary sparse graph and the additional grid lines, the OARSMT-guided sparse graph is generated. An example is shown in Figure 7(c). When performing maze routing on the generated graph, a valid solution without any obstacle violations can be obtained, as shown by the solid line example in Figure 7(c). Note that the solution on the OARSMT-guided sparse graph might not be obtainable on an ordinary sparse graph such as Figure 7(a). Besides, each net’s additional grid lines are only applied for its own maze routing, so each OARSMT-guided sparse graph has a limited complexity increase compared to the ordinary one. As a result, the proposed method has comparable execution time with better quality than that of the ordinary graph.
(2) Obstacle-aware Sparse Maze Routing: After utilizing OARSMT-guided sparse maze routing, the overflow and obstacle-violated nets experience a significant reduction; however, there are still some nets with obstacle violations. As shown in the dotted OARSMT in Figure 7(e), the OARSMT algorithm considers only obstacles within the bounding box during the tree generation stage, which may result in invalid solutions. To fix this problem, the obstacle-aware sparse maze routing can be applied exclusively to the remaining nets with violations. Utilizing the additional grid lines extending from the boundaries of all obstacles and pins for maze routing, by combining these lines with the ordinary sparse graph, an obstacle-aware sparse graph can be created, as shown in Figure 7(e) and (f). The solid lines illustrate a solution for obstacle-aware sparse maze routing. This method ensures that most nets successfully navigate around obstacles. Although the presence of numerous obstacles can slow down the obstacle-aware sparse maze routing process, it is only applied to a limited number of remaining nets. As a result, the overall runtime for this step remains reasonable.
@@ 6.1 Experiment Setup
The experiments are conducted on a server with Intel Xeon Gold 6348 CPU. In the experiments of the OARSMT algorithm, there are two types of benchmarks: Randomized testcases and standard benchmarks. Each testcase in benchmarks contains one net to be tested, which follows the regular evaluation process as previous OARSMT works [11, 4, 21]. For randomized testcase generation, 3 factors are set: The number of pins and obstacles, and the obstacle density on the layout. Other parameters are randomly generated. Each setting is tested 50 times to obtain the average result. Besides, we also test all the standard benchmarks, IND, RC, and RT [22] involved in the compared works [11, 4, 21]. $k_{l}=5$ and $k_{m}=2$ in the experiments. For OARSMT comparison, the previously advanced work proposed by Lin et al. [11], a recent work [21] and a traditional work [4] are involved, with works [11, 4] being reproduced. Similar to recent work [12], our experiments also apply the algorithm in work [11] by 2D mode. The 2D mode of work [11] does not omit any key techniques from the 3D mode, and thus it still maintains the quality with faster speed. Besides, as the tree generation in regular global routing flow typically generate 2D trees, applying work [11] under 2D mode enables the comparison under a complete global routing flow.
In the experiments of obstacle-avoiding global routing, the official open-source implementation of global router CUGR2.0 [20, 23] is adopted, where the proposed algorithms are implemented to replace the existing routing phases of CUGR2.0. However, our experimental results for wirelength, via count, and overflow cost were obtained by the evaluation function of the official CUGR2.0 implementation without change. Furthermore, to better demonstrate the effectiveness of our method, obstacle violations are evaluated separately and not included in the overflow cost metric.
ISPD24 contest [24] benchmark circuits were used in our experiments, as they contain testcases from the regular size to the very large size. We set the macros in the benchmarks as obstacles for our experiments, of which the areas are forbidden to be crossed by wires at all layers. In addition, we also add 60 random obstacles to each testcase of the benchmarks as additional testcases.
@@ 6.2 OARSMT Comparisons on Randomized Testcases
Table 1: The comparisons with the previously advanced work [11] on the randomized testcases.
@@
Wirelength Difference (<<0 Stands for Improvements)
@@
The OARSMT algorithm is first tested and compared. During the OARSMT algorithm comparisons, all obstacles in the layout are considered, rather than only considering those in the net’s bounding box as the initial routing flow. In this case, there is no violation in the experiments of the OARSMT algorithm itself.
The wirelength difference and speedup of the proposed algorithm are shown in Table 1, compared to previously advanced work [11]. In Table 1, each row represents a setting of the number of pins and each column represents a setting of the number of obstacles. Each group contains four runtime results corresponding to different obstacle densities. One can notice that the proposed algorithm has $\sim$10x - 2700x speedup from small cases to large cases.
The absolute values of the runtime for the proposed algorithm are also tested, and they are 0.63 ms for small cases and only 0.19 s for large cases. This not only shows its high speed, but also proves the scalability.
Besides, Table 1 also shows the wirelength difference, and negative results mean that the proposed algorithm has a smaller wirelength. The wirelength difference ranges from -5.76% to 2.68%. It is shown that the proposed algorithm can also outperform work [11] in the wirelength for various cases, and the wirelength improvement is 1.14% on average. For relatively large designs, the proposed algorithm shows >>3% wirelength improvement. The experiments indicate that the proposed algorithm has better wirelength for randomized testcases compared with the previous work, but also achieves orders of magnitude shorter runtime.
[…]
Finally, we briefly report the wirelength and runtime comparisons with previous 2D OARSMT work [4]. The speedup of the proposed work compared to work [4] ranges from $\sim$11x - 7374x, with 1306x on average. Meanwhile, The wirelength difference ranges from -10.04% to -0.47%. The average wirelength improvement among all cases is 5.03%. The previous work [4] results in a worse wirelength because it trims some optimized Steiner nodes when building an initial tree structure at the early stages. This again indicates the advantages of the proposed algorithm on both quality and speed.
@@ 6.3 OARSMT Comparisons on Standard Benchmarks
cases #Pin/#Obs Wirelength Runtime (s)
@@
For OARSMT algorithms, besides the comprehensive randomized testcases and comparisons, we also compare the proposed algorithm on all standard benchmarks [22] adopted in the compared works, including the previously advanced work proposed by Lin et al. [11], the recent work GSLS [21], and the traditional work [4]. Each testcase generates one OARSMT for one net within a layout with obstacles. The results are shown in Table 6.3. The speedup of our work ranges from $\sim$150x to 5800x, with limited wirelength overhead traded compared to the works [11, 21], and with even better wirelength compared to the work [4]. Note that for Lin et al. [11], there is no result for RC12 in either the reproduction or the original report, due to the large runtime. In contrast, the proposed algorithm only needs $\sim$0.2s for RC12. Besides, the runtime of the proposed algorithm also shows the scalability as it increases almost linearly with increasing problem size, and this is in contrast with recent work GSLS [21].
Figure 9: Performance comparison: our OARSMT vs. Lin et al.’s method with decreased iterations.
Furthermore, we conducted an additional experiment to compare Lin et al.’s method, as shown in Figure 9, based on the same industrial benchmark in Table 6.3. The method proposed by Lin et al. can balance runtime and wirelength by adjusting the optimization. By default, the iteration ends when the cost does not decrease after one iteration. To ensure fair comparison and better demonstrate the superiority of our algorithm, we try to decrease its iteration to get better runtime by gradually sacrificing the wirelength, even when Lin et al.’s wirelength is already adjusted to be worse than our proposed OARSMT method, Lin et al.’s runtime is still much higher throughout the testcases. This indicates a better trade-off of our proposed rule-based OARSMT algorithm.
In complex designs with large numbers of nets, using the previous OARSMT algorithm [4, 11] during the tree generation stage may become so computationally expensive that it exceeds the runtime of all other steps combined in a traditional routing flow. In conclusion, the experiments on standard benchmarks also demonstrate the high efficiency and scalability of the proposed algorithm. This is not only because of the theoretical lower time complexity, but also because of the simpler processing.
@@ 6.4 Proposed Obstacle-Avoiding Global Routing Evaluations
  • • The Lin et al. is our proposed obstacle-avoiding global routing flow with the previously advanced OARSMT [11].
@@
As shown in Table 3, we compare the proposed global router (“Ours”) with CUGR 2.0 [20], enhanced CUGR 2.0, and our global routing flow integrated with the previously advanced OARSMT [11] method (“Lin et al.”). The enhanced CUGR 2.0 is CUGR 2.0 with the increased edge cost for obstacles during maze routing. The cost is large enough to effectively make the maze routing avoid the obstacles in the sparse graphs if possible. We separately compare wirelength, via count, obstacle violation cost, overflow cost, and runtime. Note that the overflow cost does not include the obstacle violation cost.
In Table 3, CUGR 2.0 performs poorly in terms of obstacle violations. Although the enhanced CUGR 2.0 has shown significant improvement, it still fails to effectively eliminate obstacle violations. In contrast, the proposed global routing flow (“Ours”) can reduce the final obstacle violation to zero, which proves its effectiveness. Additionally, “Ours” demonstrates reductions in wirelength and overflow in most test cases when compared to both the original and enhanced CUGR 2.0, achieving an average improvement of 1.96% in wirelength and a 28.06% reduction in overflow cost compared to enhanced CUGR 2.0, although with acceptable overhead in terms of via count and runtime, as more detours are required to navigate around obstacles. The effectiveness of obstacle elimination and overflow is primarily due to the challenges posed by complex obstacle distributions. In traditional algorithms, treating obstacles similarly to overflow complicates the balancing of costs in each stage. This disrupts the collaboration of stages in a traditional global routing process. Meanwhile, maze routing sparse grids in traditional global routers without OARSMT guidance also limit the opportunities of avoiding obstacles. In this case, addressing obstacle violations directly in the early stage and modifying the later stages to incorporate early guidance allows global routing stages to effectively manage obstacles. This is also supported by the insights presented in Section 6.5.2.
To further evaluate the effectiveness of the proposed complete flow, we also replace the OARSMT algorithm in the proposed flow with the previously advanced OARSMT [11] (“Lin et al.” in Table 3). It shows that although both “Ours” and “Lin et al.” eliminate all obstacle violations in the benchmark, “Lin et al.” spends 9.21x runtime on average than “Ours”, with similar wirelength, via count, and overflow cost. This further demonstrates the efficiency of our obstacle-avoiding global routing flow with the proposed OARSMT algorithm.
Figure 10: Comparison of CUGR 2.0, CUGR 2.0 enhanced, and proposed method across different stages in Wirelength, Via Count, Overflow, Runtime and Obstacle Violations.
In addition, we conducted an experiment to analyze the performance of each stage. The Figure 10 illustrates a comparison between CUGR 2.0, enhanced CUGR 2.0, and our method ("Ours") in three stages: initial routing (IR), initial routing with congestion augmentation (IRCA), and rip-up and reroute (RRR). The comparison is based on the average values of all cases for five metrics: wirelength, via count, overflow, runtime, and obstacle violations. As shown in Figure 10, our initial routing exhibits a higher overflow due to our priority consideration of avoiding obstacles. Although traditional global routing has less overflow in the early stages, they do not effectively consider obstacles. This is why, in Figure 10, during initial routing, traditional routers have better wirelength and overflow, but they have worse obstacle violations. In later stages, without considering the obstacles in advance or the guidance for rerouting, transitional routers have higher difficulties handling the violations, leading to even worse wirelength and overflow, as well as remaining violations.
6.5 Obstacle-Avoiding Global Routing Insight Analysis
In addition to the results of the complete flow, we also analyze the insides, including runtime breakdown, the comparisons for the initial routing step, the comparisons of using different graphs during maze routing, the effectiveness of OARSMT-guided sparse maze routing, and ablation experiments isolating rule-based OARSMT and proposed maze routing separately.